OpenStudy (anonymous):

2x+8=14 what is the value of x in thos equation

OpenStudy (anonymous):

a. x=3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

b. x=6 c.x=7 d.x=11 r u sure

OpenStudy (anonymous):

no

OpenStudy (anonymous):

... The question is are YOU sure. It doesn't matter if he can do it or not. He's not the one being tested/graded.

OpenStudy (anonymous):

yea u tell her polpak

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh and btw tht is the right answer

OpenStudy (anonymous):

Or is it?

OpenStudy (anonymous):

i just need help :(

OpenStudy (anonymous):

yes it is ! if u dont trust me then what is your answer oh and cece21 it is x=3

OpenStudy (anonymous):

I told you how to do this the last time. Subtract 8 from both sides. What do you have left?

OpenStudy (anonymous):

one i have zero and the other side i have 14-8 = 6

OpenStudy (anonymous):

sabb: As I said before... Trust isn't the issue. The issue is whether or not cerce has convinced herself that you are correct. If not then she's just going to have to ask you how to do the next problem, and the one after that, etc. And finally she will do badly on the test when you're not there to do it for her.

OpenStudy (anonymous):

polpak hows tht

OpenStudy (anonymous):

No, you don't have 0. You still have 2x. And I showed you step by step how to do it.

OpenStudy (anonymous):

Go back and look at the example I gave again, and do this one the same way I did that one.

OpenStudy (anonymous):

Polpak this is a website were people give each other a answer and i was doing the right thing by giving her the answer

OpenStudy (anonymous):

2x+8=14 -8-8 0 6 8 divided by 0 is 0 and 6 divied by 8 =0.75

OpenStudy (anonymous):

6x2 = 12 0r 6-2x=4x

OpenStudy (anonymous):

Sabb, this is not a website for giving answers. It's a website for getting help. They aren't the same thing. Clearly cece has no idea how you arrived at your solution, and unless you want to just do all her homework and tests for her you aren't really helping.

OpenStudy (anonymous):

polpak thts not important just help me i have a bid test toww

OpenStudy (anonymous):

Cece \(2x+8=14\) -8 -8 ======= 2x = 6

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

You can't just ignore the 2x. x is what you are trying to find.

OpenStudy (anonymous):

is it going be 2x0=6

OpenStudy (anonymous):

And as I said, giving you the answer (as sabb already did nearly 15 mins ago hasn't helped you any)

OpenStudy (anonymous):

2x = 6

OpenStudy (anonymous):

i kno im trying to see the information though

OpenStudy (anonymous):

Ok look at it this way.

OpenStudy (anonymous):

when u do 2x is the mulipty

OpenStudy (anonymous):

6 2x-?

OpenStudy (anonymous):

no. x is the variable. it is not multiplication.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

you need to review your basics. You have clearly missed some of the fundamentals before getting to this material.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can u help me with the basic

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think i hav the answer now

OpenStudy (anonymous):

helllp

OpenStudy (anonymous):

if you have the answer, what do you need help with?

OpenStudy (anonymous):

r u gonna check it

OpenStudy (anonymous):

Not if you don't show it..

OpenStudy (anonymous):

2x+8=14 -8 8 you cross out 8 then u go to the right side and - 14-8 = 6 2x 2 cross out the 2's and leave the x and you do 6 divided by 2 which = 3

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can u help me with this question